OpenStudy (anonymous):

9/27 as a fraction in simpiliest form

OpenStudy (anonymous):

Welcome to Open Study :)

OpenStudy (anonymous):

\[\frac{ 9 }{ 27 }\] Factor out the GCF from the numerator and denominator.

OpenStudy (goformit100):

First of all @abcd123456 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

what do the 9 (numerator) and the 27 (denominator) have in common? what factor?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

hmm that could work, but it is a longer way. That is why i asked for the greatest, yes 3 is a factor, but what is the greatest common factor?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

This is when you take out the factor of 3. \[\frac{ 9 \div 3 }{ 27 \div 3 } = \frac{ 3 }{ 9 }\] As you see, that is still not the simplest/reduced form.

OpenStudy (anonymous):

soo show me plz

OpenStudy (anonymous):

Can you try it?

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

what exactly do you not understand?

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

So you said to take out a factor of 3 from both the numerator and denominator and you get the following: \(\frac{ 9 \div 3 }{ 27 \div 3 } = \frac{ 3 }{ 9 }\) But \[\frac{ 3 }{ 9 }\] is still not the simplest, most reduced form. So you still need to factor out something else. (what do the 3 (numerator) and the 9 (denominator) have in common? what factor?

OpenStudy (anonymous):

You could have just factored out the GCF, which in this case is 9. This would have made it more easier and simpler + get you to the answer faster. 3 is a factor of 9 and 27 but it is not the GCF (Greatest common factor). 9 is the GCF. So it would be like the following: \(\frac{ 9 }{ 27 } \rightarrow \frac{ 9 \div 9}{ 27 \div 9 } \rightarrow ?\)